# Quantum Polymorphisms and Commutativity Gadgets

### v1

@@ 5. Intermezzo: Contextuality and bifurcations
Let $\mathbb{X},\mathbb{A}$ be digraphs, let $Q\colon\mathbb{X}\to_{q}\mathbb{A}$ be a quantum homomorphism, and let $\omega$ be a positive integer. A bifurcation for $Q$ is a set
$V=\{(x_{0},a_{0}),(x_{0},a_{0}^{\prime})\}\cup\{(x_{1},a_{1}),\dots,(x_{\omega-1},a_{\omega-1})\}\cup\{(x_{\omega},a_{\omega}),(x_{\omega},a_{\omega}^{\prime})\}\subseteq X\times A$
such that
  • $x_{0},x_{1},\dots,x_{\omega}$ is a path in $\mathbb{X}$;
@@
  • $Q_{\alpha,\beta}Q_{\delta,\gamma}\neq 0$ for each $((\alpha,\beta),(\delta,\gamma))\in V^{2}\setminus\{((x_{0},a_{0}^{\prime}),(x_{\omega},a_{\omega}^{\prime})),((x_{\omega},a_{\omega}^{\prime}),(x_{0},a_{0}^{\prime}))\}$ such that $\alpha\neq\delta$. We call $\omega$ the length of the bifurcation.
Recall that $\mathbb{X}^{u}$ denotes the undirected graph underlying a digraph $\mathbb{X}$.
Theorem 26.
Let $\mathbb{X}$ and $\mathbb{A}$ be digraphs such that $\mathbb{X}^{u}$ is connected, and let $Q\colon\mathbb{X}\to_{q}\mathbb{A}$ be a contextual quantum homomorphism. Then there exists a bifurcation for $Q$ whose length is at most the diameter of $\mathbb{X}^{u}$.
We will need the following result from [manvcinska2016quantum], of which we give an alternative proof making use of the machinary of Section 3.
Proposition 27 ([manvcinska2016quantum]).
Let $Q\colon\mathbb{X}\to_{q}\mathbb{Y}$ be a quantum homomorphism between two digraphs $\mathbb{X}$ and $\mathbb{Y}$. If $x,x^{\prime}\in X$ and $y,y^{\prime}\in Y$ are such that there exists a directed $\ell$-walk from $x$ to $x^{\prime}$ in $\mathbb{X}$ but there is no directed $\ell$-walk from $y$ to $y^{\prime}$ in $\mathbb{Y}$, then $Q_{x,y}Q_{x^{\prime},y^{\prime}}=0$.
@@
Consider the structure $\mathbb{G}$ consisting of a path of length $\ell$, and call $u,v$ the two extremities of this path. Since this structure is a tree, Proposition 47 implies that $(\mathbb{G},u,v)$ is a q-definition for certain binary relations $R_{X}\subseteq X^{2},R_{Y}\subseteq Y^{2}$ in $\mathbb{X}$ and $\mathbb{Y}$, respectively. It follows from Proposition 17 that $Q$ is a quantum homomorphism $(X,R_{X})\to_{q}(Y,R_{Y})$. By assumption, $R_{X}$ contains $(x,x^{\prime})$ and $R_{Y}$ does not contain $(y,y^{\prime})$, and therefore $Q_{x,y}Q_{x^{\prime},y^{\prime}}=0$ by $($$\mathbf{QH_{1}}$)). ∎
We shall repeatedly use a simple fact about projectors, stated below without proof.
Fact 28.
Let $P,Q$ be projectors onto some Hilbert space $H$. Then $[P,Q]=0$ if and only if $PQP=PQ=QP$.
Proof of Theorem 26.
Let $Q\colon\mathbb{X}\to_{q}\mathbb{A}$ be a contextual quantum homomorphism over some Hilbert space $H$ of dimension $h$. Let $S$ be the set of pairs $(x,x^{\prime})\in X^{2}$ for which there exist $(a,a^{\prime})\in A^{2}$ such that $Q_{x,a},Q_{x^{\prime},a^{\prime}}]\neq 0$. Since $Q$ is contextual, $S$ is non-empty. Choose $(x,x^{\prime})\in S$ in a way that $\operatorname{dist}(x,x^{\prime})$ is minimum over all pairs in $S$ (where the distance is computed in $\mathbb{X}^{u}$), and call such distance $\omega$. Note that $\omega\geq 2$, as $Q$ satisfies $($$\mathbf{QH_{2}}$)). Let $x=x_{0},x_{1},\dots,x_{\omega}=x^{\prime}$ be a shortest oriented path in $\mathbb{X}$ connecting $x$ and $x^{\prime}$. By the minimality of $\operatorname{dist}(x,x^{\prime})$, observe that the PVMs $Q_{x_{\alpha}}$ and $Q_{x_{\beta}}$ commute for each $\alpha\leq\beta\in\{0,\dots,\omega\}$ such that $\alpha\neq 0$ or $\beta\neq\omega$. In particular, we can choose an o.n. basis $\mathcal{E}=(e_{1},\dots,e_{h})$ for $H$ that diagonalises all PVMs $Q_{x_{\alpha}}$ for each $\alpha\in\{0,\dots,\omega-1\}$. On the other hand, since $Q_{x_{0}}$ and $Q_{x_{\omega}}$ do not commute, there exist $a_{0},a_{\omega}\in A$ such that $Q_{x_{0},a_{0}},Q_{x_{\omega},a_{\omega}}]\neq 0$.
Take the set $Z=\{i\in[h]:Q_{x_{\omega},b}e_{i}\in\{0,e_{i}\}\}$, and consider the projectors $P_{1}$ onto the space $\operatorname{span}(\{e_{i}:i\in Z\})$ and $P_{2}$ onto the space $\operatorname{span}(\{e_{i}:i\in[h]\setminus Z\})$. Observe that $P_{1}P_{2}=0$ and $P_{1}+P_{2}=\operatorname{id}$. Moreover, we have
(4) $[P_{1},M]=[P_{2},M]=0\quad\forall M\in\{Q_{x_{\alpha},d}:\alpha\in\{0,\dots,\omega-1\},d\in A\}\cup\{Q_{x_{\omega},a_{\omega}}\}.$
@@
since, for $i\in Z$, $e_{i}$ is an eigenvector of $Q_{x_{\omega},a_{\omega}}$. Hence, $P_{1}$ also commutes with $Q_{x_{\omega},a_{\omega}}$. Finally, $P_{1},M]+[P_{2},M]=[\operatorname{id},M]=0$ for each $M$ because of the bilinearity of the commutator, so eq. 4 holds. In particular, the operators $Q_{x_{0},a_{0}}P_{1}$, $Q_{x_{0},a_{0}}P_{2}$, $Q_{x_{\omega},a_{\omega}}P_{1}$ and $Q_{x_{\omega},a_{\omega}}P_{2}$ are all projectors. Observe now that
@@
(5) $+[Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{2}].$
Combining eq. 4 and 28, we find that the second summand in the right-hand side of section 5 is
@@
since $P_{1}P_{2}=0$. Similarly, we obtain $Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{1}]=0$. Consider now the first summand in the right-hand side of section 5. Both the projectors $Q_{x_{0},a_{0}}P_{1}$ and $Q_{x_{\omega},a_{\omega}}P_{1}$ are diagonal in the basis $\mathcal{E}$. Hence, they commute, too. We deduce that
(6) $[Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{2}]=[Q_{x_{0},a_{0}},Q_{x_{\omega},a_{\omega}}]\neq 0.$
We now claim that the matrices in the set $\{Q_{x_{\alpha},d}P_{2}:\alpha\in[\omega-1],d\in A\}\cup\{Q_{x_{\omega},a_{\omega}}P_{2}\}$ pairwise commute. The fact that $Q_{x_{\alpha},d}P_{2},Q_{x_{\alpha}^{\prime},d^{\prime}}P_{2}]=0$ for each $\alpha,\alpha^{\prime}\in[\omega-1],d,d^{\prime}\in A$ is trivial, since $\mathcal{E}$ diagonalises both. Moreover,
@@
where the first and third equalities come from eq. 4 and 28, while the second holds since the PVMs $Q_{x_{\alpha}}$ and $Q_{x_{\omega}}$ are assumed to commute by the minimality of $\operatorname{dist}(x,x^{\prime})$. So, the claim is true, and we can simultaneously diagonalise the matrices in the set $\{Q_{x_{\alpha},d}P_{2}:\alpha\in[\omega-1],d\in A\}\cup\{Q_{x_{\omega},a_{\omega}}P_{2}\}$ in some o.n. basis $\mathcal{F}=(f_{1},\dots,f_{h})$ of $H$.
For the next step, it will be useful to define the sets $L,M\subseteq[h]$ such that $Q_{x_{0},a_{0}}P_{2}=\sum_{\ell\in L}e_{\ell}e_{\ell}^{*}$ and $Q_{x_{\omega},a_{\omega}}P_{2}=\sum_{m\in M}f_{m}f_{m}^{*}$. For $m\in M$, consider also the set $A_{m}=\{i\in[h]:{\langle e_{i},f_{m}\rangle}\neq 0\}$. Observe that, for any $m\in M$,
@@
(7) $\xi_{i}\in[h]\setminus Z\quad\mbox{ for both }i=1\mbox{ and }i=2.$
@@
Fix now some $\alpha\in[\omega-1]$, and recall that the PVM $Q_{x_{\alpha}}$ is diagonalised by $\mathcal{E}$. Hence, there exists exactly one element $c_{1}\in A$ such that $Q_{x_{\alpha},c_{1}}e_{\xi_{1}}=e_{\xi_{1}}$, and exactly one element $c_{2}\in A$ such that $Q_{x_{\alpha},c_{2}}e_{\xi_{2}}=e_{\xi_{2}}$. We claim that $c_{1}=c_{2}$. Suppose this is not the case. It is time to remember that we chose $\mathcal{F}$ in such a way that it diagonalises both $Q_{x_{\omega},a_{\omega}}P_{2}$ and all the matrices $Q_{x_{\alpha},d}P_{2}$ for $d\in A$. Hence, we can write $Q_{x_{\alpha},c_{1}}P_{2}=\sum_{p\in V_{1}}f_{p}f_{p}^{*}$ and $Q_{x_{\alpha},c_{2}}P_{2}=\sum_{p\in V_{2}}f_{p}f_{p}^{*}$ for some $V_{1},V_{2}\subseteq[h]$. Since $Q_{x_{\alpha}}$ is a PVM, we have $Q_{x_{\alpha},c_{1}}Q_{x_{\alpha},c_{2}}=0$. Using eq. 4 and 28, we deduce that
@@
Observe now that
@@
since $\xi_{2}\in A_{\bar{m}}$. Hence, $Q_{x_{0},a_{0}^{\prime}}Q_{x_{\omega},a_{\omega}}\neq 0$. Note also that the $Q_{x_{0},a_{0}}Q_{x_{\omega},a_{\omega}}=0$ is trivial from the fact that $Q_{x_{0},a_{0}},Q_{x_{\omega},a_{\omega}}]\neq 0$.
To conclude, we claim that there exists some $a_{\omega}^{\prime}\in A$ such that $a_{\omega}^{\prime}\neq a_{\omega}$ and
(9) $\left(\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}\right)Q_{x_{\omega},a_{\omega}^{\prime}}\neq 0.$
Note that this would be enough to finish the proof. Indeed, it follows from the minimality of $\operatorname{dist}(x,x^{\prime})$ that (9) guarantees that $Q_{x_{\omega},a_{\omega}^{\prime}}$ commutes with $Q_{x_{i},a_{i}}$ for each $0\leq i\leq\omega-1$.
Suppose, for the sake of contradiction, that the claim is false. Then, the fact that $Q_{x_{\omega}}$ is a PVM would imply that
@@
a contradiction. Hence, the claim is true.
Putting all together, we conclude that the vertices
$\{(x_{0},a_{0}),(x_{0},a_{0}^{\prime})\}\cup\{(x_{1},a_{1}),\dots,(x_{\omega-1},a_{\omega-1})\}\cup\{(x_{\omega},a_{\omega}),(x_{\omega},a_{\omega}^{\prime})\}$
form a bifurcation for $Q$ of length $\omega$, and the proof is complete. ∎
@@
Figure 2. A bifurcation of length $\omega$. Solid lines mean non-orthogonality.
6. Quantum polymorphisms of odd cycles
The goal of this section is to prove the following.

### v2

@@ 7 Intermezzo: Contextuality and bifurcations
Let $\mathbb{X},\mathbb{A}$ be relational structures, let $Q\colon\mathbb{X}\to_{\operatorname{q}}\mathbb{A}$ be a quantum homomorphism, and let $\omega$ be a positive integer. Take a $4$-tuple $\mathcal{B}=(\mathbf{x},\mathbf{a},a_{0}^{\prime},a_{\omega}^{\prime})$, where
  • $\mathbf{x}=(x_{0},\dots,x_{\omega})\in X^{\omega+1}$,
  • $\mathbf{a}=(a_{0},\dots,a_{\omega})\in A^{\omega+1}$, and
  • $a_{0}^{\prime},a_{\omega}^{\prime}\in A$.
We say that $\mathcal{B}$ is a (contextuality) bifurcation for $Q$ if it meets the following five conditions:
  1. $(i)$
    $x_{0},x_{1},\dots,x_{\omega}$ is a walk in $\operatorname{Gaif}(\mathbb{X})$;
  2. $(ii)$
@@
  3. $(iii)$
    $Q_{x_{i},a_{i}}Q_{x_{j},a_{j}}\neq 0$ for each $i<j\in\{0,\dots,\omega\}$;
  4. $(iv)$
    $Q_{x_{0},a_{0}^{\prime}}Q_{x_{j},a_{j}}\neq 0$ for each $j\in\{1,\dots,\omega\}$;
  5. $(v)$
    $Q_{x_{i},a_{i}}Q_{x_{\omega},a_{\omega}^{\prime}}\neq 0$ for each $i\in\{0,\dots,\omega-1\}$.
We call $\omega$ the length of the bifurcation $\mathcal{B}$.
Figure 3 and Figure 4 illustrate two examples of contextuality bifurcations. The following is the main result of this section, showing that bifurcations occur in any contextual quantum homomorphism, provided that the left structure is connected.
See 8 We shall derive Theorem 8 as a consequence of the next, slightly stronger result.
Proposition 66.
Let $\mathbb{X}$ and $\mathbb{A}$ be relational structures and let $Q\colon\mathbb{X}\to_{\operatorname{q}}\mathbb{A}$ be a quantum homomorphism. Let $x_{0},x_{1},\dots,x_{\omega}$ be a walk in $\operatorname{Gaif}(\mathbb{X})$, and choose two values $a,\tilde{a}\in A$ in a way that
  • $(I)$
    $Q_{x_{0},a},Q_{x_{\omega},\tilde{a}}]\neq 0$, but
  • $(II)$
    for all $i,j\in\{0,\dots,\omega\}$ such that $\{i,j\}\neq\{0,\omega\}$, the PVMs $Q_{x_{i}}$ and $Q_{x_{j}}$ commute.
Then $Q$ has a bifurcation $\mathcal{B}=(\mathbf{x},\mathbf{a},a_{0}^{\prime},a_{\omega}^{\prime})$ for some tuple $\mathbf{a}=(a_{0},\dots,a_{\omega})\in A^{\omega+1}$ such that $a_{0}=a$, $a_{\omega}=\tilde{a}$ and two values $a_{0}^{\prime},a_{\omega}^{\prime}\in A$, where $\mathbf{x}=(x_{0},\dots,x_{\omega})$.
In the proof of Proposition 66, we shall repeatedly use the following fact about projectors.
Fact 67.
Let $P,Q$ be projectors onto some Hilbert space $H$. Then $[P,Q]=0$ if, and only if, $PQP=PQ=QP$.
@@ 5. $(v)$
If $[P,Q]=0$, then $PQ=QP=QP^{2}=PQP$. The other direction is clear. ∎
[…]
Figure 4: A bifurcation of length $\omega$. Solid lines mean non-orthogonality.
Proof of Proposition 66.
Let $Q\colon\mathbb{X}\to_{\operatorname{q}}\mathbb{A}$ be a quantum homomorphism over some Hilbert space $H$ of dimension $h$, and consider a walk $x_{0},x_{1},\dots,x_{\omega}$ in $\operatorname{Gaif}(\mathbb{X})$ and values $a,\tilde{a}\in A$ that satisfy conditions $(I)$ and $(II)$ of the proposition. Note that condition $(I)$ forces $\omega\geq 2$, as $Q$ satisfies $($$\mathbf{QH_{2}}$)). Moreover, using condition $(II)$, we can choose an orthonormal basis $\mathcal{E}=(e_{1},\dots,e_{h})$ for $H$ that diagonalises all PVMs $Q_{x_{\alpha}}$ for each $\alpha\in\{0,\dots,\omega-1\}$.
Take the set $Z=\{i\in[h]:Q_{x_{\omega},{\tilde{a}}}e_{i}\in\{0,e_{i}\}\}$, and consider the following endomorphisms of $H$: the projector $P_{1}$ onto the space $\operatorname{span}(\{e_{i}:i\in Z\})$, and the projector $P_{2}$ onto the space $\operatorname{span}(\{e_{i}:i\in[h]\setminus Z\})$. Observe that $P_{1}P_{2}=0$ and $P_{1}+P_{2}=\operatorname{id}$. Moreover, calling $a_{0}=a$ and $a_{\omega}=\tilde{a}$, we have
$[P_{1},M]=[P_{2},M]=0\quad\forall M\in\{Q_{x_{\alpha},d}:\alpha\in\{0,\dots,\omega-1\},d\in A\}\cup\{Q_{x_{\omega},a_{\omega}}\}.$ (10)
@@
since, for each $i\in Z$, $e_{i}$ is an eigenvector of $Q_{x_{\omega},a_{\omega}}$. Hence, $P_{1}$ also commutes with $Q_{x_{\omega},a_{\omega}}$. Finally, $P_{1},M]+[P_{2},M]=[\operatorname{id},M]=0$ for each $M$ because of the bilinearity of the commutator, so eq. 10 holds. In particular, the operators $Q_{x_{0},a_{0}}P_{1}$, $Q_{x_{0},a_{0}}P_{2}$, $Q_{x_{\omega},a_{\omega}}P_{1}$, and $Q_{x_{\omega},a_{\omega}}P_{2}$ are all projectors. Observe now that
@@
$+[Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{2}].$ (11)
Combining eq. 10 and 67, we find that the second summand in the right-hand side of section 7 is
@@
since $P_{1}P_{2}=0$. Similarly, we obtain $Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{1}]=0$. Consider now the first summand in the right-hand side of section 7. Both projectors $Q_{x_{0},a_{0}}P_{1}$ and $Q_{x_{\omega},a_{\omega}}P_{1}$ are diagonal in the basis $\mathcal{E}$. Hence, they commute, too. We deduce that
$[Q_{x_{0},a_{0}}P_{2},Q_{x_{\omega},a_{\omega}}P_{2}]=[Q_{x_{0},a_{0}},Q_{x_{\omega},a_{\omega}}]\neq 0.$ (12)
We now claim that the matrices in the set $\{Q_{x_{\alpha},d}P_{2}:\alpha\in[\omega-1],d\in A\}\cup\{Q_{x_{\omega},a_{\omega}}P_{2}\}$ pairwise commute. The fact that $Q_{x_{\alpha},d}P_{2},Q_{x_{\alpha}^{\prime},d^{\prime}}P_{2}]=0$ for each $\alpha,\alpha^{\prime}\in[\omega-1],d,d^{\prime}\in A$ is clear, since $\mathcal{E}$ diagonalises both. Moreover,
@@
where the first and third equalities come from eq. 10 and 67, while the second holds since the PVMs $Q_{x_{\alpha}}$ and $Q_{x_{\omega}}$ are assumed to commute by condition $(II)$. So, the claim is true, and we can simultaneously diagonalise the matrices in the set $\{Q_{x_{\alpha},d}P_{2}:\alpha\in[\omega-1],d\in A\}\cup\{Q_{x_{\omega},a_{\omega}}P_{2}\}$ in some orthonormal basis $\mathcal{F}=(f_{1},\dots,f_{h})$ of $H$.
For the next step, it is useful to define the sets $L,M\subseteq[h]$ such that $Q_{x_{0},a_{0}}P_{2}=\sum_{\ell\in L}e_{\ell}e_{\ell}^{*}$ and $Q_{x_{\omega},a_{\omega}}P_{2}=\sum_{m\in M}f_{m}f_{m}^{*}$. For $m\in M$, consider also the set $A_{m}=\{i\in[h]:{\langle e_{i},f_{m}\rangle}\neq 0\}$. Observe that, for any $m\in M$,
@@
$\xi_{i}\in[h]\setminus Z\quad\mbox{ for both }i=1\mbox{ and }i=2.$ (13)
@@
Fix now some $\alpha\in[\omega-1]$, and recall that the PVM $Q_{x_{\alpha}}$ is diagonalised by $\mathcal{E}$. Hence, there exists exactly one element $c_{1}\in A$ such that $Q_{x_{\alpha},c_{1}}e_{\xi_{1}}=e_{\xi_{1}}$, and exactly one element $c_{2}\in A$ such that $Q_{x_{\alpha},c_{2}}e_{\xi_{2}}=e_{\xi_{2}}$. We claim that $c_{1}=c_{2}$. Suppose this is not the case. It is time to remember that we chose $\mathcal{F}$ in such a way that it diagonalises both $Q_{x_{\omega},a_{\omega}}P_{2}$ and all the matrices $Q_{x_{\alpha},d}P_{2}$ for $d\in A$. Hence, we can write $Q_{x_{\alpha},c_{1}}P_{2}=\sum_{p\in V_{1}}f_{p}f_{p}^{*}$ and $Q_{x_{\alpha},c_{2}}P_{2}=\sum_{p\in V_{2}}f_{p}f_{p}^{*}$ for some $V_{1},V_{2}\subseteq[h]$. Since $Q_{x_{\alpha}}$ is a PVM, we have $Q_{x_{\alpha},c_{1}}Q_{x_{\alpha},c_{2}}=0$. Using eq. 10 and 67, we deduce that
@@
We now make the following claim.
Claim 68.
There exists some $a_{\omega}^{\prime}\in A$ such that $a_{\omega}^{\prime}\neq a_{\omega}$ and
$\left(\prod_{0\leq i\leq\omega-1}Q_{x_{i},a_{i}}\right)Q_{x_{\omega},a_{\omega}^{\prime}}\neq 0.$ (15)
The truth of the claim would be enough to conclude the proof of the proposition. Indeed, consider the 4-tuple $(\mathbf{x},\mathbf{a},a_{0}^{\prime},a_{\omega}^{\prime})$, where $\mathbf{x}=(x_{0},\dots,x_{\omega})$ and $\mathbf{a}=(a_{0},\dots,a_{\omega})$.
Observe first that
@@
since $\xi_{2}\in A_{\bar{m}}$. Hence, $Q_{x_{0},a_{0}^{\prime}}Q_{x_{\omega},a_{\omega}}\neq 0$. Note also that $Q_{x_{0},a_{0}}Q_{x_{\omega},a_{\omega}}\neq 0$, as we know that $Q_{x_{0},a_{0}},Q_{x_{\omega},a_{\omega}}]\neq 0$.
All in all, the above argument proves conditions $(iii)$ and $(iv)$ of Definition 65, while condition $(i)$ is clear from the initial assumptions and $(ii)$ follows from the choice of $a_{0}^{\prime}$ and from 68. Moreover, $(v)$ also follows from 68 by noting that condition $(II)$ of the current proposition guarantees that $Q_{x_{\omega},a_{\omega}^{\prime}}$ commutes with $Q_{x_{i},a_{i}}$ for each $1\leq i\leq\omega-1$ and, thus, in order for (15) to hold it must be the case that $Q_{x_{i},a_{i}}Q_{x_{\omega},a_{\omega}^{\prime}}\neq 0$ for each $i\in\{0,\dots,\omega-1\}$. As a consequence, $\mathcal{B}=(\mathbf{x},\mathbf{a},a_{0}^{\prime},a_{\omega}^{\prime})$ is a bifurcation for $Q$, as required.
We are left to prove 68. Suppose, for the sake of contradiction, that the claim is false. Then, the fact that $Q_{x_{\omega}}$ is a PVM would imply that
@@
a contradiction. Hence, 68 is true, and the proof of Proposition 66 is complete. ∎
Having established Proposition 66, we can now prove the main result of this section.
Proof of Theorem 8.
@@
Remark 69.
Observe that the proofs of both Proposition 66 and Theorem 8 do not make use of the fact that the given quantum homomorphism $Q\colon\mathbb{X}\to_{\operatorname{q}}\mathbb{A}$ satisfies condition $($$\mathbf{QH_{1}}$)). Hence, both statements hold for any quantum function $Q\colon X\to_{\operatorname{q}}A$ satisfying $($$\mathbf{QH_{2}}$)).
As a simple application of Theorem 8, we can prove the following non-contextuality result.
Proposition 70.
Let $\mathbb{X},\mathbb{Y}$ be relational structures such that $\operatorname{Gaif}(\mathbb{X})$ and $\operatorname{Gaif}(\mathbb{Y})$ are odd cycles of the same length. Then every quantum homomorphism $\mathbb{X}\to_{\operatorname{q}}\mathbb{Y}$ is non-contextual.
In particular, it follows that quantum endomorphisms of odd cycles are non-contextual.
Corollary 71.
For each odd $m\geq 3$, every quantum homomorphism $Q\colon\mathbb{C}_{m}\to_{\operatorname{q}}\mathbb{C}_{m}$ is non-contextual.
We point out that a different proof of Corollary 71 was obtained in [34] via a result from [4] on the quantum symmetry of vertex-transitive graphs.
Proof of Proposition 70.
Suppose first that $\mathbb{X}$ is not essentially binary—i.e., recalling Definition 15, $\mathbb{X}$ contains some relation $R^{\mathbb{X}}$ such that at least one tuple $\mathbf{x}\in R^{\mathbb{X}}$ contains at least three distinct entries—call them $x,y,z$. Then the set $\{x,y,z\}\subseteq X$ induces a $\mathbb{K}_{3}$ in $\operatorname{Gaif}(\mathbb{X})$, and the assumption on $\operatorname{Gaif}(\mathbb{X})$ implies that $X$ cannot contain any other vertex. Hence, in this case, the fact that each quantum homomorphism from $\mathbb{X}$ to $\mathbb{Y}$ is non-contextual immediately follows from $($$\mathbf{QH_{2}}$)).
Suppose now that $\mathbb{X}$ is essentially binary and assume, for the sake of contradiction, that $\mathbb{X},\mathbb{Y}$ admit some contextual quantum homomorphism $Q\colon\mathbb{X}\to_{\operatorname{q}}\mathbb{Y}$. Let $m$ be the number of vertices in $\operatorname{Gaif}(\mathbb{X})=\operatorname{Gaif}(\mathbb{Y})$. Applying Theorem 8, we deduce that $Q$ has a bifurcation $\mathcal{B}=(\mathbf{x},\mathbf{a},a_{0}^{\prime},a_{\omega}^{\prime})$ of length $\omega\leq\frac{m-1}{2}$. Let as usual $\mathbf{x}=(x_{0},\dots,x_{\omega})$ and $\mathbf{a}=(a_{0},\dots,a_{\omega})$, and consider the pairs $(x_{0},a_{0}),(x_{0},a_{0}^{\prime}),(x_{1},a_{1}),(x_{2},a_{2})\in X\times Y$. By Definition 65, we know that $Q_{x_{0},a_{0}}Q_{x_{1},a_{1}}\neq 0$, $Q_{x_{0},a^{\prime}_{0}}Q_{x_{1},a_{1}}\neq 0$, $Q_{x_{1},a_{1}}Q_{x_{2},a_{2}}\neq 0$, $Q_{x_{0},a_{0}}Q_{x_{2},a_{2}}\neq 0$, and $Q_{x_{0},a_{0}^{\prime}}Q_{x_{2},a_{2}}\neq 0$. Using Proposition 16, we deduce that $\{a_{0},a_{1}\},\{a_{0}^{\prime},a_{1}\}$, and $\{a_{1},a_{2}\}$ are all edges of $\operatorname{Gaif}(\mathbb{Y})$. Note also that $x_{0}$ and $x_{2}$ lie on an (m2)(m-2)-walk in $\operatorname{Gaif}(\mathbb{X})$. Since no walk of such length connects a vertex to itself in $\operatorname{Gaif}(\mathbb{Y})$Proposition 16 ensures that $a_{0}\neq a_{2}$ and $a_{0}^{\prime}\neq a_{2}$. Finally $a_{0}\neq a_{0}^{\prime}$ again by Definition 65. We deduce that $a_{1}$ has degree at least three in $\operatorname{Gaif}(\mathbb{Y})$, a contradiction. ∎
8 Quantum polymorphisms of the Siggers digraph
In this section, we investigate the complexity of the quantum CSP parameterised by the Siggers digraph—i.e., the digraph obtained by taking a directed triangle and making one of its directed edges undirected, see Figure 5. The significance of this directed graph is that it is the smallest digraph whose classical CSP is NP-complete.
021
Figure 5: The Siggers digraph.
We prove the following result.
See 10
Throughout this section, we denote the Siggers digraph by $\mathbb{A}$ (and, as usual, we denote its vertex set $\{0,1,2\}$ by $A$). In order to deduce the undecidability of $\operatorname{CSP_{q}}(\mathbb{A})$ from the algebraic machinery developed in the previous sections, we show that all quantum polymorphisms of the Siggers digraph are non-contextual by making use of contextuality bifurcations.